# Quantised Global Autoencoder: A Holistic Approach to Representing Visual Data

### v1

@@ E Alternative Formulations
Staying with the ImageNet example, even when applying a few layers of convolution and pooling first, when combining e.g. $16$-by-$16$ feature dimensions with $256$ channels into $64\cdot 1024$ (to have enough space for quantisation and to produce $1024$ tokens), this again would require over $4$ billion parameters. Again, this is not feasible for both hardware requirements and generalisation.
@@ G Additional Results, Qualitative Analysis, Comparison Against VQ-VAE
While we did not extensive train and fine-tune our approach on ImageNet[5] due to a lack of benchmarks and resources (e.g. no comparison to VQ-WAE[33] possible due to undisclosed code), we do provide some results on (unoptimised) settings to demonstrate that we can also perform well on larger datasets and with larger numbers of tokens. We train on ImageNet in $64$-by-$64$ pixels with $256$ and $1024$ tokens. We also train our approach on $128$-by-$128$ pixels with $256$ tokens, shown in Fig. 14.
We show further results in ImageNet in $128$ pixels in Fig. 14, comparing against an already improved VQ-VAE with codebook regularisation that avoids unused codebook entries. In result, our approach is better at producing sharper features, even if they are at times slightly misplaced.
H Generative Performance of VQ-VAE versus QG-VAE Tokens
While generation is not our focus, and these are only early results without fine tuning or architecture optimisation, we still provide a small quantitative measurement of applying our learned representation in generation.
We also measured the performance of applying a generative transformer[13] on tokens generated by our approach. We compare to a VQ-VAE[31], while for reasons of fairness, we capped both at the same reconstruction quality, and applied the transformer only in unconditional form. We show the resulting FID plots in Fig. 11, and additional randomly selected result in Fig. 12.
While the initial improvements are smaller and FID is higher, we argue that this is a result of our more complex latent space: Simply repeating a token e.g. for blue sky is not beneficial for our representation. However, with the later better results, we argue that this shows how our learned, natively linear sequence of tokens is beneficial for generation tasks when compared to artificially cutting a $2$-dimensional grid of tokens into strips. We argue that further research in this direction, in particular on larger datasets, could be promising.
Figure 11: Generation quality when training the same transformer on our pseudo frequency tokens of CIFAR-10[15] versus training on tokens produced by a VQ-VAE[31]. Our approach performs better, producing lower FID. We compute on $10,000$ unconditional samples each.
@@ H Generative Performance of VQ-VAE versus QG-VAE Tokens
Input VQ-VAE QG-VAE (ours) Input VQ-VAE QG-VAE (ours)
Input VQ-VAE QG-VAE (ours) Input VQ-VAE QG-VAE (ours)
Figure 14: Examples on ImageNet $128$, always input, output (VQ-VAE[31]), output (QG-VAE). Note how small details, like the spider legs or global structures like the fence behind the bike, are better represented for our approach instead of just being blurred. Each image has only $256$ tokens ($288$ bytes).

### v2

@@ E Alternative Formulations
Staying with the ImageNet example, even when applying a few layers of convolution and pooling first, when combining e.g. $16$-by-$16$ feature dimensions with $256$ channels into $64\cdot 1024$ (to have enough space for quantisation and to produce $1024$ tokens), this again would require over $4$ billion parameters. Again, this is not feasible in the sense of both hardware requirements and generalisation.
@@ G Additional Results, Qualitative Analysis, Comparison Against VQ-VAE
While we did not extensive train and fine-tune our approach on ImageNet[5] due to a lack of benchmarks and resources (e.g. no comparison to VQ-WAE[33] possible due to undisclosed code), we do provide some results on (unoptimised) settings to demonstrate that we can also perform well on larger datasets and with larger numbers of tokens. We train on ImageNet in $64$-by-$64$ pixels with $256$ and $1024$ tokens. We also train our approach on $128$-by-$128$ pixels with $256$ tokens, shown in Fig. 15, comparing against an already improved VQ-VAE with codebook regularisation that avoids unused codebook entries. In result, our approach is better at producing sharper features, even if they are at times slightly misplaced. We also give additional examples produced by QGGAN, i.e. a QG-VAE with additional VQGAN[7]-like sharpening.
H Generative Results
We provide additional randomly selected result of random autoregressive generation of our QGVAE in Fig. 13.
@@ H Generative Results
Input VQ-VAE QG-VAE (ours) QGGAN (ours)
Input VQ-VAE QG-VAE (ours) QGGAN (ours)
Figure 15: Examples on ImageNet $128$, always input, output (VQ-VAE[31]), output (QG-VAE), output (QGGAN). Note how small details, like the spider legs or global structures like the fence behind the bike, are better represented for our approach instead of just being blurred. Each image has only $256$ tokens ($288$ bytes).